On monday, Hank drove to work at an average speed of 70km/h and arrived 1 minute late. On Tuesday, he left at the same time and took the same route. This time he drove at an average speed of 75km/h and arrived 1 minute early. How long is his route to work?
Answer by edjones(8007)

st=d
s*t=s*t
70*60(t+1)=75*60(t-1) Making speed into km/min is easier.
4200t+4200=4500t-4500
-300t=-8700
t=29 min
70 * 30/60 = 35 km
.
Ed